OpenStudy (anonymous):

Implicit differentiation. I'm getting stuck on how to solve this problem x^2y+y^2x=-2

OpenStudy (luigi0210):

Did you at least try differentiating it yet?

OpenStudy (anonymous):

yeah, i used product rule but i couldn't factor out the dy/dx properly

OpenStudy (luigi0210):

If I differentiated it properly it's \((2xy+x^2y')+(2yy'x+y^2)=0\) Whoops, kind of already did it for him >.< But am I right?

OpenStudy (anonymous):

Yeah that's how far I got

OpenStudy (luigi0210):

Move anything that doesn't have a y' to right

OpenStudy (luigi0210):

*to the right

OpenStudy (anonymous):

So that would have the y^2 and 2xy on the right side of the equation and then when you factor out the y' on the left, divide?

OpenStudy (luigi0210):

Yup: \((2xy+x^2y')+(2yy'x+y^2)=0\) \(x^2y'+2yy'x=-y^2-2xy\) \(y'(x^2+2y)=-y^2-2xy\) And just divide and you're done.

OpenStudy (luigi0210):

Should be 2yx not 2y, my mistake.

OpenStudy (anonymous):

Thank you so much for that, I've been stuck on this for a while now. :)

OpenStudy (luigi0210):

You're welcome, and good luck :)